OpenStudy (freemap):

What are the values of x when tan(x) is undefined? A.-270°, -90°, 90°, and 270° B.-270°, -90°, 0°, 90°, and 270° C.-360°, -180°, 0°, 180°, and 360° D.-360°, 0° and 360°

OpenStudy (anonymous):

remember tan= sin/cos

OpenStudy (anonymous):

where does cos=0?

OpenStudy (freemap):

I'm sorry can you give me some background information of tan and sin and cos

OpenStudy (anonymous):

so tangent is the same as sin/cos

OpenStudy (anonymous):

anything divided by 0 is undefined

OpenStudy (anonymous):

so using the unit circle, find where cos=0

OpenStudy (freemap):

Isn't undefined a vertical line?

OpenStudy (anonymous):

undefined means that nothing exists at that point

OpenStudy (anonymous):

So if you look at the unit circle cos is 0 at 90 degrees

OpenStudy (anonymous):

can you find the other point that it equals 0?

OpenStudy (freemap):

270 degrees

OpenStudy (freemap):

Is the answer b

OpenStudy (anonymous):

270 is correct. What does cos equal at 0 degrees?

OpenStudy (freemap):

I don't know 0

OpenStudy (anonymous):

no

OpenStudy (freemap):

270 and 90 degrees

OpenStudy (anonymous):

at 0 cos=1

OpenStudy (anonymous):

that means that cos is defined at 0

OpenStudy (anonymous):

so now we have our two values 90 and 270

OpenStudy (anonymous):

The negative values will also produce 0 for cos

OpenStudy (anonymous):

So what is the answer?

OpenStudy (freemap):

A

OpenStudy (jdoe0001):

\(\bf tan(\theta)=\cfrac{sin(\theta)}{{\color{brown}{ cos(\theta) }} }\qquad if\ {\color{brown}{ cos(\theta) }}=0\ then \\ \quad \\ tan(\theta)=\cfrac{sin(\theta)}{{\color{brown}{ 0}}}\impliedby undefined\) as already pointed out

OpenStudy (anonymous):

That's right

OpenStudy (freemap):

SO the answer is A

OpenStudy (jdoe0001):

and at 90 and 270 degrees cosine is indeed 0 going backwards with negatives, will land on the same spots as 10115658 said

OpenStudy (anonymous):

http://intmstat.com/trigonometric-functions/tan2.gif

OpenStudy (anonymous):

looking at the graph of tan, the lines disappear at 90, 270, -90, -270

OpenStudy (freemap):

Thanks so much for your time.

OpenStudy (anonymous):

no problem